Staff:

Lbto: Juan Carlos

Inaf: Felice, eta l (telescope control room)

Forecast.

20230121_forecast0.png

20230121_forecast1.png

Details.

4:35PM

Power on SoulSXWFS.

OCAM ok

04:56PM
Steve called the night.
"Still ice on the top of the shutter, it is not safe to open it"

6:18PM

Running a close loop checkout.

Deploy the argos cal unit.

setup luci.


Sun Jan 22 01:21:57.546 2023 aos.PresetAO{FLAO}.started - PresetAO {FLAO}[2753] (aomode:ACE-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:0[0],soy:0,rox:-0.01[-0.01],roy:-0.03,mag:10,cindex:0,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)
Sun Jan 22 01:21:57.563 2023 aos.info - WFS source set to: FLAOWFS
disturbToAdd string "/local/aomeas/adsec_calib/CMD/disturb/dist_flao2_KL_v20_atm_s1.6_L040.0_v15.0_ovfreq1700.00_sd3892.fits"


optmize

20230122_013758

1.14,0.36,0.41

gopton

TN:+luci

_014330,_014541

psf oscillating at low frequencies.20mas total jitter

optimize:20230122_014907

modalSR 65

_015413

new preset.



optimize:_020550

1.14,0.4,0.6

still a tilt oscillation

30mas oscillation.

_021052

Mirror ventillation in ON.

7:26PM

StopAO

PArk ACU

Adsec rest.

WFs power off.

-- JuanCarlosGuerra - 21 Jan 2023
Topic revision: r2 - 22 Jan 2023, JuanCarlosGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback